# SpotFormer: Multi-Scale Spatio-Temporal Transformer for Facial Expression Spotting

### v2

@@ 5 Conclusion
In this paper, we proposed an efficient framework for FES. First, we proposed an SW-MRO feature to magnify motion information while alleviating head movement issues. Such an optical flow feature is designed to perceive complete MEs and discern the differences between MaEs and MEs. Then, we presented SpotFormer, a powerful graph-based Transformer designed to learn multi-scale spatio-temporal relationships from SW-MRO features for frame-level probability estimation. In SpotFormer, we proposed an FLGP operation for downsampling facial graph-structured data to achieve multi-scale spatial feature fusion. Moreover, learnable convolution layers are employed for multi-scale temporal feature learning. In addition, we introduced supervised contrastive learning into SpotFormer to enhance the discriminability between different types of expressions at the frame level. Extensive comparative experiments and ablation studies on the SAMM-LV, CAS(ME)2, and CAS(ME)3 datasets demonstrated the effectiveness of our proposed framework, particularly in ME spotting.
Our model also has some limitations. Performance under conditions with fewer available MEs requires improvement, as evidenced by the experimental results on the CAS(ME)2 dataset. Although our proposed SW-MRO feature is effective in extracting robust motion features, it is not computationally friendly and highly relies on the accuracy of current optical flow extraction method. Therefore, our method may become less effective under extreme lighting conditions. In future work, we aim to explore an efficient method for solving the data-imbalance problem existing in current datasets and develop an end-to-end spot-then-recognize framework for facial expression analysis.
@@ Acknowledgments
[Uncaptioned image] Yicheng Deng received the B.Eng. and M.Eng. degrees from Beijing Jiaotong University, Beijing, China, in 2019 and 2022, respectively. He is now pursuing his Ph.D. degree in Osaka University. His research interests include facial expression analysis and video understanding.
[Uncaptioned image] Hideaki Hayashi (S’ 13–M’ 16) received the B.E., M.Eng, and D.Eng. degrees from Hiroshima University, Hiroshima, Japan, in 2012, 2014, and 2016 respectively. He was a Research Fellow of the Japan Society for the Promotion of Science from 2015 to 2017 and an assistant professor in Department of Advanced Information Technology, Kyushu University from 2017 to 2022. He is currently an associate professor with the D3 Center, Osaka University. His research interests focus on neural networks, machine learning, and medical data analysis.
[Uncaptioned image] Hajime Nagahara is a professor at the D3 Center, Osaka University, since 2017. He received Ph.D. degree in system engineering from Osaka University in 2001. He was a research associate of the Japan Society for the Promotion of Science from 2001 to 2003. He was an assistant professor at the Graduate School of Engineering Science, Osaka University, Japan from 2003 to 2010. He was an associate professor in Faculty of information science and electrical engineering at Kyushu University from 2010 to 2017. He was a visiting associate professor at CREA University of Picardie Jules Verns, France, in 2005. He was a visiting researcher at Columbia University in 2007-2008 and 2016-2017. Computational photography and computer vision are his research areas. He received an ACM VRST2003 Honorable Mention Award in 2003, IPSJ Nagao Special Researcher Award in 2012, ICCP2016 Best Paper Runners-up, and SSII Takagi Award in 2016. He is a program chair in ICCP2019, Associate Editor for IEEE Transaction on Computational Imaging in 2019-2022, and Director of Information Processing Society of Japan in 2022-2024.

### v3

@@ 5 Conclusion
In this paper, we proposed an efficient framework for FES. First, we proposed an CSW-MRO feature to magnify motion information while alleviating head movement issues. Such an optical flow feature is designed to perceive complete MEs and discern the differences between MaEs and MEs. Then, we presented SpotFormer, a powerful graph-based Transformer designed to learn multi-scale spatio-temporal relationships from CSW-MRO features for frame-level probability estimation. In SpotFormer, we proposed an FLGP operation for downsampling facial graph-structured data to achieve multi-scale spatial feature fusion. Moreover, learnable convolution layers are employed for multi-scale temporal feature learning. In addition, we introduced supervised contrastive learning into SpotFormer to enhance the discriminability between different types of expressions at the frame level. Extensive comparative experiments and ablation studies on the SAMM-LV, CAS(ME)2, and CAS(ME)3 datasets demonstrated the effectiveness of our proposed framework, particularly in ME spotting.
Our model also has some limitations. Performance under conditions with fewer available MEs requires improvement, as evidenced by the experimental results on the CAS(ME)2 dataset. Although our proposed CSW-MRO feature is effective in extracting robust motion features, it is not computationally friendly and highly relies on the accuracy of current optical flow extraction method. Therefore, our method may become less effective under extreme lighting conditions. In future work, we aim to explore an efficient method for solving the data-imbalance problem existing in current datasets and develop an end-to-end spot-then-recognize framework for facial expression analysis.